OpenStudy (anonymous):

find the second derivative of the function: y=(2-2x^2+x^6)/x^9

OpenStudy (anonymous):

\[y'=\frac{(-4^x+6x^5)}{x^9}-\frac{(9*(2-2x^2+x^6))}{x^{10}}\]

OpenStudy (anonymous):

Is that the first or second derivative? (d^2y)/(dx^2)?

OpenStudy (anonymous):

\[y''=\frac{(-4+30x^4)}{x^9}-\frac{(18(-4x+6x^5))}{x^{10}}+\frac{90(2-2x^2+x^6)}{x^{11}}\]

OpenStudy (anonymous):

there is another way to solve this is by dividing all the terms by x^9

OpenStudy (anonymous):

do you how to find the derivative of this form- numerator/denominator ,where both numerator and denominator are functions of some variable??

OpenStudy (anonymous):

Can you please explain, Siddharth18?

OpenStudy (anonymous):

y=(2-2x^2+x^6)/x^9 we can rewrite this now \[y = 2x^{-9}-2x^{-7}+x^{-3}\]

OpenStudy (anonymous):

if u differentiate this function it would be much easier than using the quotient rule, that i just used earlier...

OpenStudy (anonymous):

well there is a standard rule to solve such a form.